# A note on the omega-chaos

### v1

@@ Abstract.
A continuous map $f\colon X\to X$ is said to be transitive if for any non-empty open subsets $U,V$ of $X$, there is $i\geq 0$ such that $f^{i}(U)\cap V\neq\emptyset$. Note that $f$ is transitive if and only if $X=\omega(x,f)$ for some $x\in X$.
@@
Let $f\colon X\to X$ be a continuous map and let $g=f^{\mathbb{N}}\colon X^{\mathbb{N}}\to X^{\mathbb{N}}$. If there is a closed subset $\Lambda$ of $X$ with $f(\Lambda)\subset\Lambda$ such that
  • $\Lambda\not\in\{\emptyset,X\}$,
  • $f|_{\Lambda}\times f\colon\Lambda\times X\to\Lambda\times X$ is transitive,
then $g$ is $\omega$-chaotic.
@@
Since $f|_{\Lambda}\times f\colon\Lambda\times X\to\Lambda\times X$ is transitive, we have
$\omega((p,z),f|_{\Lambda}\times f)=\Lambda\times X$
for some $(p,z)\in\Lambda\times X$. It follows that
@@
Note that
  • $X=\omega(z,f)$,
  • $X$ is not a finite set.
Since $X(=\omega(z,f))$ is perfect, i.e., has no isolated point, $X\setminus\Lambda$ is perfect and locally compact Hausdorff; therefore, $X\setminus\Lambda(=\omega(z,f)\setminus\Lambda)$ is uncountable. Since
@@
by Theorem 1, we conclude that $g$ is $\omega$-chaotic. ∎
@@
By Remark 1 and [1, Theorem 1], we obtain the following lemma.
@@
Let $f\colon X\to X$ be a continuous map. A non-empty closed subset $M$ of $X$ with $f(M)\subset M$ is called a minimal set for $f$ if $M=\omega(x,f)$ for all $x\in X$, or equivalently, for any closed subset $S$ of $M$, $f(S)\subset S$ implies $S\in\{\emptyset,M\}$. An $x\in X$ is called a minimal point for $f$ if
  • $x\in\omega(x,f)$,
  • $\omega(x,f)$ is a minimal set for $f$.
Let $M(f)$ denote the set of minimal points for $f$ and note that $M(f)$ is the (disjoint) union of minimal sets for $f$. By Zorn’s lemma, there is at least one minimal set for $f$; therefore, $M(f)\neq\emptyset$.
@@
  • as a consequence of Furstenberg’s $2$ implies $n$ theorem (see [3, Proposition II.3]), if $f$ is weakly mixing, then so is $g$,

### v2

@@ Abstract.
A continuous map $f\colon X\to X$ is said to be transitive if for any non-empty open subsets $U,V$ of $X$, there is $i\geq 0$ such that $f^{i}(U)\cap V\neq\emptyset$. Note that $f$ is transitive if and only if $X=\omega(x,f)$ for some $x\in X$. A non-empty closed subset $M$ of $X$ with $f(M)\subset M$ is called a minimal set for $f$ if $M=\omega(x,f)$ for all $x\in M$. We say that $f$ is non-minimal if $X$ is not a minimal set for $f$.
@@
Let $f\colon X\to X$ be a continuous map and let $g=f^{\mathbb{N}}\colon X^{\mathbb{N}}\to X^{\mathbb{N}}$. If $f$ is transitive and non-minimal, then $g$ is $\omega$-chaotic.
@@
Since $f$ is transitive, we have $X=\omega(z,f)$ for some $z\in X$. By the Auslander–Ellis theorem (see, e.g., [1, Theorem 5]), there are a minimal set $M$ for $f$ and $p\in M$ such that
$\liminf_{i\to\infty}d(f^{i}(p),f^{i}(z))=0$
and so
@@
Note that $X\setminus M\neq\emptyset$ because $f$ is non-minimal. Since $X(=\omega(z,f))$ is perfect, i.e., has no isolated point, $X\setminus M$ is perfect and locally compact Hausdorff, thus $X\setminus M(=\omega(z,f)\setminus M)$ is uncountable. Since
@@
letting $\Lambda=M$, we see that the assumptions of Theorem 1 are satisfied; therefore, $g$ is $\omega$-chaotic. ∎
@@
By Remark 1 and [2, Theorem 1], we obtain the following lemma.
@@
Given a continuous map $f\colon X\to X$, we denote by $M(f)$ the (disjoint) union of minimal sets for $f$. A non-empty closed subset $M$ of $X$ with $f(M)\subset M$ is a minimal set for $f$ if and only if for any closed subset $S$ of $M$, $f(S)\subset S$ implies $S\in\{\emptyset,M\}$. By Zorn’s lemma, there is at least one minimal set for $f$; therefore, $M(f)\neq\emptyset$.
@@
  • as a consequence of Furstenberg’s $2$ implies $n$ theorem (see [4, Proposition II.3]), if $f$ is weakly mixing, then so is $g$,